OpenStudy (anonymous):

Solve the inequality. Type your answer using interval notation. 4|2x+7|-20≤0

OpenStudy (anonymous):

since its an absolute value, dont i have to do something with 2x+7<= to 5 and -5?

OpenStudy (anonymous):

Would it be [-6,-1]?

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

\[|2x+7|\leq 5\] \[-5\leq 2x+7\leq 5\] \[-12\leq 2x\leq -2\] \[-6\leq x\leq -1\]

OpenStudy (anonymous):

so yes, it would be \[[-6,-1]\]

OpenStudy (anonymous):

Thanks